OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1)

hartnn (hartnn):

subtract 2nd equation from 1st equation. tell me what u get.

mathslover (mathslover):

\[\large{x-2y=14}\] \[\large{x+3y=9}\] solve for x in the first equation

OpenStudy (anonymous):

it cant be a or d

mathslover (mathslover):

x = 14+2y now put this value in second one: \[\large{(14+2y)+3y=9}\]

hartnn (hartnn):

thats correct @Carol329

mathslover (mathslover):

Can you solve for y now? @Carol329

OpenStudy (anonymous):

its c :)

mathslover (mathslover):

correct

hartnn (hartnn):

wow! how do u figure that out ?

mathslover (mathslover):

good work @Carol329 .